OpenStudy (anonymous):

If (x+a) is a factor of 2Xsquare(two X square)+2ax+5x+10,find a ? guys i need the procedure;

OpenStudy (lgbasallote):

\[\large 2x^2 + 2ax + 5x + 10\] that's the polynomial right?

OpenStudy (anonymous):

2?

OpenStudy (lgbasallote):

@gadivamsi is there an = 0? or that's really just the question??

OpenStudy (anonymous):

ya lgbasallote,it is 2x2+2ax+5x+10

OpenStudy (lgbasallote):

but is there an = 0?

OpenStudy (anonymous):

u kno that's an expression....nd it's got a factor

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

hmm im gonna assume that there's = 0 there... well if x+a is a factor that means x = -a is a solution if x = -a is a solution then i can substitute -a into every x \[2(-a)^2 + 2a(-a) + 5a + 10 = 0\] \[2a^2 - 2a^2 + 5a + 10 = 0\] \[5a + 10 = 0\] \[5a = -10\] you can solve for a easily now...

OpenStudy (lgbasallote):

im just thinking how you can solve for a \(\textbf{IF}\) there's no = 0

OpenStudy (anonymous):

expression's dont have =0's u kno............wen they have a factor say x+a.......then.............f(-a) wud be 0 :|

OpenStudy (anonymous):

what you did is right but the answer should be +2 not -2

OpenStudy (lgbasallote):

GREAT! im right then!

OpenStudy (anonymous):

it's 2

OpenStudy (lgbasallote):

\[5a = -10\] divide both sides by 5 \[a = -2\]

OpenStudy (anonymous):

hmm i knew this but the answer in 2

OpenStudy (anonymous):

u kno it's 5 * -a

OpenStudy (lgbasallote):

ohh yeag

OpenStudy (lgbasallote):

yeah*

OpenStudy (lgbasallote):

i overlooked that lol \[-5a = -10\] divide both sides by -5 \[a = 2\] lol there you go

OpenStudy (lgbasallote):

so i guess there was an implicit = 0 there